# T-Explainer: A Model-Agnostic Explainability Framework Based on Gradients

### v2

@@ 5.2 Synthetic Data
We generated two different synthetic datasets comprising $1,000$ instances each. The first is a 4-dimensional (4-FT) dataset where each instance $\mathbf{x}$ is generated as follows. We distribute the target label $\mathbf{y}\in[0,1]$ equally across each dataset half; thus, each class has $500$ instances. Conditioned to the value of $\mathbf{y}$, we sample the instance $\mathbf{x}$ as $x_{1:2}\sim\mathcal{N}(\mu_{y},\,\Sigma_{y})$. We choose $\mu_{0}=[0,0]^{\mathrm{T}}$ and $\mu_{1}=[-2,2]^{\mathrm{T}}$, $\Sigma_{0}=[[1,1],[-1,1]]^{\mathrm{T}}$ and $\Sigma_{1}=\mathbf{I}$, where $\mu_{0}$, $\Sigma_{0}$ and $\mu_{1}$, $\Sigma_{1}$ denote the means and covariance matrices of the Gaussian distributions associated with instances in classes 00 and $1$ respectively. The features $x_{1:2}$ are called $core\_1$ and $core\_2$. Random values are assigned to the features $x_{3:4}$ ($noise\_1$ and $noise\_2$). Such configuration results in a dataset holding two predictive (important) and two random noise (non-important) features with a small mixture area between the important features to introduce some degree of complexity.
@@
The second synthetic dataset is more robust. It is a 20-dimensional (20-FT) dataset created using the OpenXAI synthetic data generation tool, whose algorithm is described in Agarwal et al. [4]. According to the authors, the algorithm ensures the creation of a dataset that encapsulates feature dependencies and clear local neighborhoods, key properties to guarantee the explanations derived from this synthetic dataset remain consistent with the behavior of the models trained on such data. Using controlled synthetic data for XAI checking meets the correctness dimension of explainability evaluation [42].
@@
We trained the XRFC and 3H-NN models for the synthetic datasets experiments. The XRFC classifier achieved $96\%$ accuracy on the 4-FT data and $83.5\%$ accuracy on the 20-FT data, while the 3H-NN model achieved $97.5\%$ accuracy on the 4-FT data and $83.5\%$ accuracy on the 20-FT synthetic dataset. Evaluating explanation methods on accurate models is good practice in controlled settings based on synthetic data [42]. The higher the model’s accuracy level, the more it can be assumed that the model adhered to the reasoning designed in the data.
@@
Note that we are not basing our benchmark experiments on synthetic datasets with massive amounts of instances. We know it is necessary to use a reasonable amount of data to train and test machine learning models. According to Aas et al. [1], well-known feature importance methods become unstable in tasks with more than ten dimensions. In this sense, data dimensionality is more critical to assessing the stability of feature importance methods than the number of instances. Thus, we generated synthetic datasets with enough instances ($1,000$ for each dataset) to train our models and assess the XAI methods.
@@
Table 1 shows the RIS, ROS, and RES stability metrics for T-Explainer, TreeSHAP, and LIME, explaining the XRFC model on the 4-FT data. Notice that TreeSHAP has the best RIS, ROS, and RES results. The good performance of TreeSHAP is expected, as the 4-FT is a low-dimensional dataset, and TreeSHAP takes advantage of it by relying on a deterministic version of Shapley values computation [7]. Table 2 shows the same metrics for the 3H-NN model on the 4-FT dataset. The T-Explainer performed considerably better than the other XAI methods in terms of RIS (7×7$\times$7 × better than DeepLIFT), ROS, and also the RES metric.
@@
Table 3 presents the PGI results of the explainers applied to the 3H-NN predictions on the 4-FT data. T-Explainer explanations are considerably more faithful than the other methods in identifying the most important feature for predictions (Top 1 column), i.e., T-Explainer’s explanations demonstrated the closest behavior to what the underlying model learned as the most important feature. When the task considered more features, the second and third most important, SHAP outperformed T-Explainer; however, our method performed consistently more faithfully to the model’s predictions than the other explainers (columns Top 2 and Top 3).
@@
Tables 4 and 5 depict the stability of the XRFC and 3H-NN models on the 20-FT dataset. From Table 4, we can see that T-Explainer is superior to TreeSHAP and LIME in the RIS and ROS metrics for both maximum and mean values. TreeSHAP reached a close performance regarding maximum ROS, but T-Explainer is considerably better on the average ROS. Similar results can be observed in Table 5, where T-Explainer outperforms all the other explainers in terms of RIS (3×3$\times$3 × better than Integrated Gradients), with slightly better performance than Integrated Gradients in the mean ROS but outperforming all the other explainers in maximum ROS.
SHAP was the most unstable method, which can be justified in the context of the 20-FT dataset because, in high-dimensional spaces, SHAP uses a random sampling algorithm rather than the deterministic one, leading to unstable explanations [7]. Moreover, SHAP is also prone to suffer from extrapolations [29, 30].
@@
In general, Tables 1, 2, 4, and 5 show that T-Explainer is robust regarding different stability metrics evaluating explanations’ consistency and continuity in controlled synthetic data tasks [42], being among the best performance methods, especially when data dimensionality is high.
Observing the PGI results in Table 6, we have T-Explainer as the most faithful explanation method to the original model predictive behavior for the most important feature (Top 1 column). Enlarging the set of the $k$ most important features (Top 3–9 columns), SHAP presented the best performances; however, T-Explainer consistently outperformed all the other explainers, being the second most faithful explanation method explaining the 3H-NN predictions on the 20-FT synthetic dataset.
@@
Table 7 demonstrates XAI explainers’ ability to approximate the original model predictions based on their feature importance attributions (see Equation 5). We used the LAP metric (considering $0.1$ and $0.05$ as tolerance errors, i.e., allowing for $10\%$ and $5\%$ of inaccuracy) to assess the Local Accuracy preservation for feature importance attributions generated to predictions of the 3H-NN model on 4-FT and 20-FT datasets. In the 4-FT columns of Table 7, T-Explainer achieved the best performances for Local Accuracy preservation, either with $0.1$ tolerance error or the more restrictive tolerance of $0.05$. Observe that all the other explainers achieved rates below $0.5$, meaning that except for T-Explainer, all the explanation methods tested violated the Local Accuracy property in more than half of their explanations (remember LAP returns frequency rates of Local Accuracy preservation).
3H-NN LAP ($\epsilon=0.1$) LAP ($\epsilon=0.05$)
XAI 4-FT 20-FT 4-FT 20-FT
T-Exp 0.557 0.494 0.490 0.311
SHAP 0.432 0.281 0.419 0.227
LIME 0.017 0.101 0.010 0.056
I-Grad 0.477 0.166 0.445 0.122
I×$\times$×Grad 0.464 0.377 0.439 0.266
DeepLIFT 0.464 0.377 0.439 0.266
Table 7: Local Accuracy preservation when explaining the 3H-NN predictions on 4-FT and 20-FT datasets.
The 20-FT columns of Table 7 show that the T-Explainer outperformed all the other feature attribution methods regarding Local Accuracy preservation on the 20-FT dataset, considering either the $0.1$ tolerance error or the tighter tolerance of $0.05$. The results in Tables 3, 6, and 7 demonstrate T-Explainer’s capacity for achieving faithful and accurate explanations, while being stable.
In summary, T-Explainer’s performance when dealing with the XRFC classifier, a non-differentiable model, was remarkable for more complex tasks (e.g., explaining predictions for the 20-FT data). Such performances have significant practical implications, suggesting that the T-Explainer might perform well even when applied to generate explanations for non-differentiable machine learning models. The following section presents evaluation results on real-world data.
@@ 5.3 Real-world Data
This Section extends the evaluations and comparisons by applying real data from different domains. Specifically, we run experiments using four well-known real-world datasets with distinct properties regarding dimensionality, the presence of categorical attributes, and size.
@@
For the experiments with Banknote Authentication and German Credit datasets, we selected the 3H-NN model. From Table 9, one can see that T-Explainer is less stable than Integrated Gradients, but it is the second-best method on average for RIS and ROS (tied with Integrated Gradients for RES).
@@
Analyzing the PGI results in Table 10, we see all the methods tied to a faithful behavior between explanations and model predictions according to the most important feature identified in their explanations. However, T-Explainer’s performance grows when considering the second and third most important features. Only SHAP and T-Explainer achieved a faithfulness higher than $0.5$ for their explanations’ three most important features. In a low-dimensional (four features, see Table 8) set such as the Banknote Authentication, it is possible to conjecture that an explainer with a PGI below $0.5$ is generally missing a highly significant feature in its explanations.
@@
Table 11 presents the stability tests for the 3H-NN model predictions on the German Credit data. The original version of the dataset has $20$ numerical and categorical attributes. Still, due to its complex categorization system, it is practically impossible to apply it to a machine-learning task directly. We pre-processed the German Credit data to clean it, resulting in a reduced version holding nine features ($4$ numerical and $5$ categorical features, see Table 8). As shown in Table 11, T-Explainer performed as the most stable explainer regarding RIS and maximum ROS perturbations, being the second best for the mean ROS.
@@
Table 12 shows close performances regarding the PGI metric among all feature importance methods for 3H-NN predictions on the German Credit data. LIME, T-Explainer, and SHAP were the most faithful explainers in the incremental verification process of top $k$ features.
@@
Tables 9 and 11 indicate that gradient-based methods are more stable to input/output perturbations than SHAP and, in some cases, also than LIME. As one can notice, T-Explainer competes quite well with other gradient-based methods regarding stable explanations. While Tables 10 and 12 place T-Explainer between the top feature importance methods for faithful explanations.
HELOC and HIGGS are more robust datasets in terms of dimensionality and size. We selected the 3H-NN and the five-layer neural network models to compare T-Explainer with the other explainers on the real-world datasets. Specifically, we trained the 3H-NN and 5H-128-NN classifiers on HELOC, achieving $71.34\%$ and $73.27\%$ accuracy, respectively. Despite the accuracy values being lower than the ones we achieved before, those performances are in line with the results reported in the literature [16].
Table 13 shows that Integrated Gradients is the most stable explainer for almost all metrics (behind LIME only for mean RIS but with very close numbers). However, we highlight the performance of T-Explainer, especially regarding the metrics’ means. Only Integrated Gradients, LIME, and T-Explainer kept the mean RIS under ten units, which can be taken as virtually the same mean stability to input perturbations. Similar results are observed in mean ROS, where Integrated Gradients, LIME, and T-Explainer achieved the smallest values, being T-Explainer the second best. Table 14 shows the explainers’ stability when applied to the five-layer neural network. T-Explainer was the most stable method, with Integrated Gradients being the second best for mean RIS and LIME as the second best for mean ROS. Only T-Explainer and Integrated Gradients achieved the highest levels of stability in RES.
@@
Tables 15 and 16 present the PGI performances of the explainers. All the methods increased their faithfulness values by including more features to the top $k$ set evaluated, indicating that the HELOC dataset has multiple features significant to the underlying model. For the 3H-NN model, the T-Explainer presented the highest PGI value considering the nine most essential features, i.e., our method could identify nine important features with high fidelity to the model predictions out of more than $20$ predictive features. For the 5H-128-NN model, T-Explainer performed better up to the three most essential features, although our method was the second with higher PGI, after Integrated Gradients, the most faithful explainer to larger top $k$ feature sets.
@@
We trained the 3H-NN and 5H-64-NN classifiers on the HIGGS data, which achieved $69.58\%$ and $66.83\%$ accuracy, respectively, close to values reported in previous works [16]. We included a set of benchmarks using a five-layer neural network here to keep some similarity with the architectures proposed in Baldi et al. [10], which explored the use of deep networks on HIGGS. Tables 17 and 18 show good stability performances of gradient-based feature importance explainers. T-Explainer outperforms them regarding RIS perturbations while being quite competitive regarding ROS and RES metrics. Such results demonstrate T-Explainer’s ability to generate consistent explanations through different levels of model complexity and data dimensionality. Note that SHAP is more unstable than LIME, with SHAP’s claimed advantageous performance predominantly seen in low-dimensional datasets [7].
@@
Tables 19 and 20 present the explainers’ PGI performances on the HIGGS dataset. LIME achieved the highest PGI values when explaining predictions made by the 3H-NN and 5H-64-NN neural networks. However, T-Explainer is the second best regarding faithful explanations when considering larger top $k$ feature sets (see Top 7 and 9 columns).
@@ 6 Computational Performance
Another challenge that XAI methods have to face is related to computational performance. Shapley-based approaches like SHAP apply sampling approximations because the exact version of the method is an NP-hard problem with exponential computing time regarding the number of features [39]. SHAP is a leading method in feature attribution tasks due to its axiomatic properties, with a range of explainers based on different algorithms, from the exact version to optimized approaches. However, according to its documentation, the most precise algorithm is only feasible for modelings that are nearly limited to $15$ features [^2] https://shap.readthedocs.io/ExactExplainer, visited on July 2024.
For a computational comparison of T-Explainer’s performance, we generated a synthetic dataset with $16$ features (16-FT) using OpenXAI, similar to the 20-FT dataset we used before (see the Synthetic Data Section). The 3H-NN model was trained with the 16-FT dataset according to the experimental setup defined in Section 5.1. We then defined ten explanation tasks, increasing the number of explanations by ten instances from the 16-FT data for each task, ranging from ten to one hundred explanations, to compare T-Explainer with three implementations of SHAP – the optimized explainer (the same method used in the stability experiments of Section 5), KernelSHAP [39], the exact explainer based on standard Shapley values computation. Figure 2 presents the computational performance of T-Explainer and SHAP.
Figure 2: Running times of T-Explainer and SHAP explainers applied to the 3H-NN model trained on 16-dimensional synthetic data.
SHAP explainer was the fastest. However, the method was the most unstable, as we demonstrated in the stability evaluations. The T-Explainer’s performance is close to KernelSHAP, while the SHAP exact explainer running time has grown exponentially. In such a performance test, we used a reduced dataset instead of the 20-FT data because applying the SHAP exact explainer in tasks with more than $16$ features is impossible. Most of the Machine Learning problems consider high-dimensional data. For example, the HIGGS is a 28-dimensional dataset with over $98$ thousand instances. With that dataset, it is impossible to apply the exact version of SHAP, and the optimized version of SHAP performs very poorly. Something similar could be said about the HELOC data, which recently motivated a challenge by FICO [25]. In this sense, T-Explainer is placed as a competitive XAI approach regarding computational resources.

### v3

@@ 5.2 Synthetic Data
We generate two different synthetic datasets comprising $1,000$ instances each. The first is a 4-dimensional (4-FT) dataset where each instance $\mathbf{x}$ is generated as follows. We distribute the target label $\mathbf{y}\in[0,1]$ equally across each dataset half; thus, each class has $500$ instances. Conditioned on the value of $\mathbf{y}$, we sample the instance $\mathbf{x}$ as $x_{1:2}\sim\mathcal{N}(\mu_{y},\,\Sigma_{y})$. We choose $\mu_{0}=[0,0]^{\mathrm{T}}$ and $\mu_{1}=[-2,2]^{\mathrm{T}}$, $\Sigma_{0}=[[1,1],[-1,1]]^{\mathrm{T}}$ and $\Sigma_{1}=\mathbf{I}$, where $\mu_{0}$, $\Sigma_{0}$ and $\mu_{1}$, $\Sigma_{1}$ denote the means and covariance matrices of the Gaussian distributions associated with instances in classes 00 and $1$, respectively. The features $x_{1:2}$ are called $core\_1$ and $core\_2$. Random values are assigned to the features $x_{3:4}$ ($noise\_1$ and $noise\_2$). Such a configuration results in a dataset holding two predictive (important) and two random noise (non-important) features with a small mixture area between the important features to introduce some degree of complexity.
@@
The second synthetic dataset is more robust. It is a 20-dimensional (20-FT) dataset created using the OpenXAI synthetic data generation tool, whose algorithm and properties are described in Agarwal et al. [3]. According to the authors, the algorithm ensures the creation of a dataset that encapsulates feature dependencies and clear local neighborhoods, key properties to guarantee the explanations derived from this synthetic dataset remain consistent with the behavior of the models trained on such data. Using controlled synthetic data for XAI checking meets the correctness dimension of explainability evaluation [33].
@@
We trained the XRFC and 3H-NN models for the synthetic datasets experiments. The XRFC classifier achieved $86.5\%$ accuracy in the 4-FT data and $83.5\%$ accuracy in the 20-FT data, while the 3H-NN model achieved $97.5\%$ accuracy in the 4-FT data and $83.5\%$ accuracy in the 20-FT synthetic dataset. Evaluating explanation methods on accurate models is good practice in controlled settings based on synthetic data [33]. The higher the model’s accuracy level, the more it can be assumed that the model adhered to the reasoning designed in the data.
@@
Note that we are not basing our benchmark experiments on synthetic datasets with massive amounts of instances. We understand that a reasonable amount of data is essential for effectively training and testing machine learning models. According to Aas et al. [1], the well-known feature importance methods become unstable in tasks with more than ten dimensions. In this sense, data dimensionality is more critical to assessing the stability of feature importance methods than the number of instances. Thus, we generate synthetic datasets with enough instances ($1,000$ for each dataset) to train our models and evaluate XAI methods.
@@
Table 1 shows the stability metrics RIS, ROS, and RES for T-Explainer, TreeSHAP, and LIME, explaining the XRFC model on the 4-FT data. Notice that TreeSHAP has the best RIS, ROS, and RES results. The good performance of TreeSHAP is expected, as 4-FT is a low-dimensional dataset, and TreeSHAP takes advantage of it by relying on a deterministic version of Shapley values computation [5]. Table 2 shows the same metrics for the 3H-NN model in the 4-FT dataset. The T-Explainer performed considerably better than the other XAI methods in RIS (7×7$\times$7 × better than DeepLIFT), ROS, and the RES metric.
@@
Table 3 presents the PGI results of the explainers applied to the 3H-NN predictions on the 4-FT data. T-Explainer explanations are considerably more faithful than the other methods in identifying the most important feature for predictions (Top 1 column), i.e., T-Explainer’s explanations demonstrated the closest behavior to what the underlying model learned as the most important feature. When the task considered more features, the second and third most important, SHAP outperformed the T-Explainer; however, our method performed consistently more faithfully to the model’s predictions than the other explainers (columns Top 2 and Top 3).
@@
Tables 4 and 5 represent the stability of the XRFC and 3H-NN models in the 20-FT dataset. In Table 4, we observe that the T-Explainer is superior to TreeSHAP and LIME in the RIS and ROS metrics for both the maximum and mean values. TreeSHAP reached a close performance regarding maximum ROS, but T-Explainer is considerably better in average ROS. Similar results can be observed in Table 5, where the T-Explainer outperforms all other explainers in terms of RIS (3×3$\times$3 × better than Integrated Gradients), with slightly better performance than Integrated Gradients in the mean ROS but outperforms the other explainers in maximum ROS.
SHAP was the most unstable method, which can be justified in the context of the 20-FT dataset because, in high-dimensional spaces, SHAP uses a random sampling algorithm rather than a deterministic one, leading to unstable explanations [5]. Moreover, SHAP is also prone to extrapolations [20, 21].
@@
In general, Tables 1, 2, 4, and 5 show that the T-Explainer is robust in different stability metrics that evaluate the consistency and continuity of explanations in controlled synthetic data tasks [33], being among the best performance methods, especially when the data dimensionality is high.
Observing the PGI results in Table 6, we have the T-Explainer as the most faithful explanation method for the predictive behavior of the original model for the most important feature (Top 1 column). Enlarging the set of the most important $k$ features (Top 3–9 columns), SHAP presented the best performances; however, T-Explainer consistently outperformed all the other explainers, being the second most faithful explanation method explaining the 3H-NN predictions on the 20-FT synthetic dataset.
@@
Figure 2 compares the RIS and ROS metrics. Once they are based on normal perturbations, we executed both metrics multiple times to evaluate the behavior of the XAI explainers and add statistical significance to our empirical experiments. We randomly selected $100$ samples from the 20-FT data (fixing the same seed throughout the experiment for reproducibility) and performed RIS and ROS metrics $10$ times each, evaluating explanations of the 3H-NN model. Similarly to previous experiments, we present the maximum and mean values of RIS and ROS over the runs. The boxplots in Figure 2 show that T-Explainer is ranked among the most stable approaches, following the earlier findings presented in Table 5.
(a) RIS.
(b) ROS.
Figure 2: Behavior of XAI methods explaining samples predicted with the 3H-NN model trained on the 20-FT dataset when running RIS and ROS metrics multiple times. The y-axes are on a logarithmic scale for better visualization.
In summary, T-Explainer’s performance when dealing with the XRFC classifier, a non-differentiable model, was remarkable for more complex tasks (e.g., explaining predictions for the 20-FT data). Such performances have significant practical implications, suggesting that the T-Explainer might perform well even when applied to generate explanations for non-differentiable machine learning models. The following section presents the evaluation results on real-world data.
@@ 5.3 Real-world Data
This Section extends the evaluations and comparisons by applying real-world data from different domains. Specifically, we run experiments using four well-known real-world datasets with distinct properties regarding dimensionality, the presence of categorical attributes, and size.
@@
For experiments with Banknote Authentication and German Credit datasets, we selected the 3H-NN model. In Table 8, one can see that T-Explainer is less stable than Integrated Gradients, but it is the second-best method on average for RIS and ROS (tied with Integrated Gradients for RES).
@@
Analyzing the PGI results in Table 9, we see that all the methods are tied to a faithful behavior between explanations and model predictions according to the most important feature identified in their explanations. However, T-Explainer’s performance grows when considering the second and third most important features. Only SHAP and T-Explainer achieved a faithfulness higher than $0.5$ for their explanations’ three most important features. In a low-dimensional (four features, see Table 7) set such as the Banknote Authentication, it is possible to conjecture that an explainer with a PGI below $0.5$ generally lacks a highly significant feature in its explanations.
@@
Table 10 presents the stability tests for the predictions of the 3H-NN model on the German Credit data. The original version of the dataset has $20$ numerical and categorical attributes. However, applying it directly to a machine learning task is difficult due to its intricate categorization system. We preprocessed the German Credit data to clean it, resulting in a reduced version holding nine features ($4$ numerical and $5$ categorical features, see Table 7). As shown in Table 10, the T-Explainer performed as the most stable explainer regarding RIS and maximum ROS perturbations, being the second best for the mean ROS.
@@
Table 11 shows close performances regarding the PGI metric among all feature importance methods for 3H-NN predictions on German Credit data. LIME, the T-Explainer, and SHAP were the most faithful explainers in the incremental verification process of the top $k$ features.
@@
Tables 8 and 10 indicate that gradient-based methods are more stable to input/output perturbations than SHAP and, in some cases, also than LIME. As one can notice, the T-Explainer competes quite well with other gradient-based methods regarding stable explanations. Although Tables 9 and 11 place the T-Explainer between the better feature importance methods for faithful explanations.
HELOC and HIGGS are more robust datasets in terms of dimensionality and size. We selected the 3H-NN and the five-layer neural network models to compare T-Explainer with the other explainers on the real-world datasets. Specifically, we trained the 3H-NN and 5H-128-NN classifiers in HELOC, achieving $71.34\%$ and $73.27\%$ of accuracy, respectively. Despite the accuracy values being lower than those we achieved before, these performances align with the results reported in the literature [12].
Table 12 shows that Integrated Gradients is the most stable explainer for almost all metrics (behind LIME only for mean RIS but with very close numbers). However, we highlight the performance of T-Explainer, especially regarding the metrics’ means. Only Integrated Gradients, LIME, and T-Explainer kept the mean RIS under ten units, which can be taken as virtually the same mean stability to input perturbations. Similar results are observed in mean ROS, where Integrated Gradients, LIME, and T-Explainer achieved the smallest values, with T-Explainer the second best. Table 13 shows the stability of the explainers when applied to the five-layer neural network. T-Explainer was the most stable method, with Integrated Gradients being the second best for mean RIS and LIME as the second best for mean ROS. Only the T-Explainer and Integrated Gradients achieved the highest levels of stability in RES.
@@
Tables 14 and 15 present the PGI performance of the explainers. All methods increased their faithfulness values by including more features in the set of top $k$ features, indicating that the HELOC dataset has multiple significant features for the underlying model. For the 3H-NN model, the T-Explainer presented the highest PGI value considering the nine most essential features, i.e., our method could identify nine important features with high fidelity to the predictions of the model out of more than $20$ predictive features. For the 5H-128-NN model, T-Explainer performed better up to the three most essential features, although our method was the second with higher PGI, after Integrated Gradients, the most faithful explainer to larger top $k$ feature sets.
@@
We trained the 3H-NN and 5H-64-NN classifiers on HIGGS data, which achieved accuracy $69.58\%$ and $66.83\%$, respectively, close to the values reported in previous work [12]. We included a set of benchmarks using a five-layer neural network here to maintain some similarity to the architectures proposed in Baldi et al. [8], which explored the use of deep networks in HIGGS. Tables 16 and 17 show good stability performances of gradient-based feature importance explainers. T-Explainer outperforms them in RIS perturbations and remains competitive in ROS and RES metrics. Such results demonstrate the ability of the T-Explainer to generate consistent explanations through different levels of model complexity and data dimensionality. Note that SHAP is more unstable than LIME, with SHAP’s claimed advantageous performance predominantly seen in low-dimensional datasets [5].
@@
Tables 18 and 19 present the PGI performance of the explainers in the HIGGS dataset. LIME achieved the highest PGI values when explaining predictions made by the 3H-NN and 5H-64-NN neural networks. However, T-Explainer is the second best in describing faithful explanations when considering larger top $k$ feature sets (see Top 7 and 9 columns).
@@ 6 Computational Performance
Another challenge that XAI methods have to face is related to computational performance. Shapley-based approaches like SHAP apply sampling approximations because the exact version of the method is an NP-hard problem with exponential computing time regarding the number of features [30]. SHAP is a leading method in feature attribution tasks due to its axiomatic properties, with a range of explainers based on different algorithms, from the exact version to optimized approaches. However, according to its documentation, the most precise algorithm is only feasible for modelings that are nearly limited to $15$ features [^4] https://shap.readthedocs.io/ExactExplainer, visited in July 2024.
For a computational comparison of the performance of the T-Explainer, we generated a synthetic dataset with $16$ features (16-FT) using OpenXAI, similar to the 20-FT dataset we used before (see the Synthetic Data Section). The 3H-NN model was trained with the 16-FT dataset according to the experimental setup defined in Section 5.1. We then defined ten explanation tasks, increasing the number of explanations by ten instances from the 16-FT data for each task, ranging from ten to one hundred explanations, to compare the T-Explainer with three implementations of SHAP – the optimized explainer (the same method used in the stability experiments of Section 5), KernelSHAP [30], the exact explainer based on standard Shapley values computation. Figure 3 presents the computational performance of T-Explainer and SHAP. The $y$-axis (time) of Figure 3 is on a logarithmic scale for better visualization purposes.
Figure 3: Running times (in log scale) of T-Explainer and SHAP explainers applied to the 3H-NN model trained on 16-dimensional synthetic data.
SHAP explainer was the fastest for the first five iterations (explaining sets with 10 to 50 instances). However, the method was the most unstable, as we demonstrated in the stability evaluations. However, the performance of the T-Explainer is quite close to that of the SHAP explainer, and it performs faster than it for sets with 80 instances and above. KernelSHAP was slightly more expensive for small sets of instances, while the SHAP exact explainer running time has grown exponentially. In such a performance test, we used a reduced dataset instead of the 20-FT data because applying the SHAP exact explainer in tasks with more than $16$ features is currently impossible. Most machine learning problems involve high-dimensional data. For example, HIGGS is a 28-dimensional dataset with over $98$ thousand instances. With that dataset, it is impossible to apply the exact version of SHAP, and the optimized version of SHAP performs very poorly. Something similar could be said about the HELOC data, which recently motivated a challenge by FICO [17]. In this sense, T-Explainer is placed as a highly competitive XAI approach regarding computational resources.